OpenStudy (anonymous):

Which function below is the inverse of f(x) = x2 − 36? x squared over 36 ±6square root of x 1 over quantity x squared minus 36 ±square root of the quantity x plus 36

OpenStudy (anonymous):

f(x) = x^2 - 36 y = x^2 - 36 .... replace f(x) with y x = y^2 - 36 .... swap x and y now solve for y to get the inverse

OpenStudy (anonymous):

ok i got A

OpenStudy (jdoe0001):

\(\bf f(x) ={\color{blue}{ y}}= {\color{brown}{ x}}^2 - 36\qquad inverse\implies {\color{brown}{ x}}= {\color{blue}{ y}}^2 - 36\) notice to get the "inverse relation", we'd firsts swap about the variables, then solve for "y"

OpenStudy (anonymous):

oh ok nevermind its C

OpenStudy (jdoe0001):

\(\bf f(x) ={\color{blue}{ y}}= {\color{brown}{ x}}^2 - 36\qquad inverse\implies {\color{brown}{ x}}= {\color{blue}{ y}}^2 - 36 \\ \quad \\ x+36=y^2\qquad \textit{then take }\sqrt{\qquad }\textit{ to both sides}\)

OpenStudy (anonymous):

wow this is confusing. so my final answer has to be D then

OpenStudy (anonymous):

consider f(x) = y y = x^2 - 36 y+36 = x^2 x = sqrt(y+36) now consider f(x) = y => x =f^-1(y) => f^-1(y) = sqrt(y+36) therefore f^-1(x) = sqrt(x+36) Did you get it now???

OpenStudy (anonymous):

yes i got it. its B lol thank you all